OpenStudy (anonymous):

What are the x-intercepts of the graph of f(x)? f(x) = -16x2 + 24x + 16

OpenStudy (muzzack):

f(x)=-16x^2 + 24x +16

OpenStudy (anonymous):

The x intercepts are: -0.5 and 2

OpenStudy (muzzack):

dont give out the answers you have to guide them

OpenStudy (anonymous):

All you have to do is graph it. https://www.desmos.com/calculator

OpenStudy (anonymous):

−16 x 16+24x+16 −256+24x+16 −240+24x 24(−10)+24x 24(−10+x)

OpenStudy (muzzack):

well you should have said that before

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

@greenlegodude thanks for that resource

OpenStudy (anonymous):

let me test it out to see how it works

OpenStudy (anonymous):

It's an amazing resource, just copy & paste your equation in the box at the left, then look at the lines crossing the x-axis. Those will be the x-intercepts.

OpenStudy (anonymous):

pretty cool now how do i figure out this part Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work.

OpenStudy (anonymous):

same equation

OpenStudy (anonymous):

The vertex is the top of the graph

OpenStudy (anonymous):

0.75, 25

OpenStudy (anonymous):

If A(-16 in the equation) is greater than 0 then it will be a minimum, but if A is less than 0, it will be a maximum. So is the graph a maximum or a minimum?

OpenStudy (anonymous):

it will be a maximum in this case

OpenStudy (anonymous):

Yes, good job.

OpenStudy (anonymous):

Btw, to find the vertex use the equation: \[-\frac{ b }{ 2a }\]

OpenStudy (anonymous):

Your equation is in the form of ax^2 + bx + c

OpenStudy (anonymous):

So your points are: \[a = -16\] \[b = 24\] \[c = 16\]

OpenStudy (anonymous):

Plug them into the equation! \[-\frac{ 24 }{ 2(-16) }\] \[-\frac{ 24 }{ 32 }\] \[0.75\]

OpenStudy (anonymous):

That's the x value of the vertex.

OpenStudy (anonymous):

You can plug in '0.75' for x in the equation and find out what 'y' is: \[f(x) = -16x^2 + 24x + 16\] \[f(x) = -16(0.75)^2 + 24(0.75) + 16\] \[f(x) = -16(0.5625) + 18.75 + 16\] \[f(x) = -9 + 18.75 + 16\] \[f(x) = 9.75 + 16\] \[f(x) = 25.75 \] So the vertex is approximately (0.75, 25.75)

OpenStudy (anonymous):

this is awesome thank you for such detailed explanation

OpenStudy (anonymous):

No problem.